Can you help me with this problem.
Find the solution sets for
(sqrt2x+1) - (sqrt x) = 1

Answer by stanbon(75887)

(sqrt2x+1) - (sqrt x) = 1
----------
Square both sides to get:
2x+1 -2sqrt(x(2x+1)) + x = 1
Simplify and isolate the remaining radical, as follows:
2sqrt(2x^2+x) =3x
Square both sides to get:
4(2x^2+x) = 9x^2
8x^2+4x = 9x^2
x^2-4x=0
x(x-4)=0
x=0 or x=4
Check the answer in the original equation:(sqrt2x+1) - (sqrt x) = 1
Does x=0? : sqrt(1)-sqrt(1)=1 : False
Does x=4? ; sqrt(9)-sqrt(4) = 1 : True
--------------
Final Answer: x=4
=====================
Cheers,
Stan H.